OpenStudy (anonymous):

Write the given sentence as an equation. Tim’s age in 7 years will be three times what it was 19 years ago.

OpenStudy (btaylor):

t+7 = 3(t-19)

OpenStudy (btaylor):

if you were to solve it: t+7=3t-57 2t=64 t=32

OpenStudy (anonymous):

your awesome

OpenStudy (anonymous):

let tim's present age be x therefore, 7 years hence his age will bi= x+7 19 years ago his age was x-19 so according to question x+7 = 3(x-19)